OpenStudy (hauntedwoodsgal):

Hi I need help with this will medal and fan im doing this for my cousin and she said I had the wrong answer so I want her to see you all are saying to correct answer so I can prove it :)

OpenStudy (hauntedwoodsgal):

Evaluate 2y + y2 when y = 6 Explain your steps using the following words: substitute, order of operations, and variable.

OpenStudy (mathmale):

Can't help but wonder why your cousin is not on OpenStudy posing these question for himself or herself. Please write "the square of y" as y^2 or \[y^2\] and not as y2. So: f(y)=2y-y^2; let y=y. Then f(6) = ???

OpenStudy (hauntedwoodsgal):

She is not allowed to she is to young 6th grade